# Optimum Spectrum Extension for PAPR Reduction of DFT-s-OFDM

### v1

@@ IV-C2 Characterizing $N_{\rm e}^{({\rm min\,PAPR})}$
As observed in Fig. 8, the PAPR behavior as a function of $N_{\rm e}$ is well characterized by the bounds (13) and (15). Accordingly, the value of $N_{\rm e}^{({\rm min\,PAPR})}$ can be closely approximated using bound (13), with slightly better precision than bound (15). For Fig. 8(e) with 16QAM, we also display the approximate bound (16) which appears to be a slightly better and simpler way to approximate $N_{\rm e}^{({\rm min\,PAPR})}$ than (13) in this case. Note, however, that the global optimum SE size obtained from simulations may correspond only to a local minimium in the PAPR bound, as shown in Fig. 8(a) – something we observed for smaller bandwidths and FDSS windows with small attenuation. A simple way to circumvent this issue is to restrict the search range for small FDSS attenuation, or inject a correction factor into the bound. Indeed, we observe that the bounds are further apart for $N_{\rm e}=0$ and become tighter as $N_{\rm e}\to N_{\rm sc}$. A practical method to compensate for this discrepancy is to approximate the PAPR as ${\rm PAPR}\;{\rm[dB]}\approx{\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)\;{\rm[dB]}-K\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)$ where $K$ is the gap between the bound and simulation at $N_{\rm e}=0$, which can empirically be evaluated.
(a) Different QAM order or bandwidth
@@
In Fig 9(b), we show that while the shift value of $L$ can impact the PAPR, the optimal SE size is nearly the same for different $L$ values. For QPSK, $N_{\rm e}^{({\rm min\,PAPR})}$ with $L$ as in (18), (19) and $L=0$ is obtained from simulations, together with the corresponding bound-based estimates. All cases are closely overlapping, and thus the bound (13) does not capture well the small variation of $N_{\rm e}^{({\rm min\,PAPR})}$ with different $L$ values. In Fig 9(b), $N_{\rm e}^{({\rm min\,PAPR})}$ is also shown for $\pi/2$-BPSK with $L$ as in (18) and $L=0$, from simulation and corresponding bounds. In this case, the range of $N_{\rm e}^{({\rm min\,PAPR})}$ is much smaller and almost constant around $5\%$, indicating that only a small amount of SE may be beneficial. Moreover, contrary to QPSK and other QAMs, as FDSS shaping increases, $N_{\rm e}^{({\rm min\,PAPR})}$ decreases.
IV-C3 Maximum PAPR gain of FDSS-SE over FDSS
Finally, Fig. 10 shows the $10^{-3}$ PAPR CCDF as a function of FDSS shaping, both with and without SE. To reach the lowest PAPR values, the shift parameter $L$ is according to (18) for $\pi/2$-BPSK and (19) for QPSK and 16-QAM; and for FDSS-SE, the corresponding optimum SE sizes $N_{\rm e}^{({\rm min\,PAPR})}$ estimated by simulation are used. For QPSK, both filter families – Hann and Kaiser windows – are considered, while for $\pi/2$-BPSK and 16-QAM only results with Hann window are shown for clarity.
For QPSK, SE with Hann window provides an almost constant PAPR gain over FDSS without SE from $1.74$ dB to $1.54$ dB. With Kaiser window, the gain of SE is the same with no FDSS but decreases slightly more to $1.31$ dB. For both window types, the lowest PAPR achieved with FDSS-SE is approximately the same for a given maximum window attenuation ripple. For 16-QAM, the PAPR gain with Hann window is also constant but smaller, at about 1.31.3dB. For $\pi/2$-BSPK, the behavior is different: the gain of SE over FDSS gradually decreases with larger shaping, from 1.51.5dB to 0.30.3dB.
@@ V-A DFT-s-OFDM Reception
where $\mathsf{snr}$ denotes the signal-to-noise ratio (SNR), $\{z[n]\}$ is an average white Gaussian noise (AWGN) with $\mathbb{E}\left[|z[n]|^{2}\right]=1$, and the channel is normalized so that $\sum_{l=0}^{P-1}\mathbb{E}\left[|h_{p}|^{2}\right]=1$.
@@ V-A1 Equalization
The combining in (24) is equivalent to [10, Eq. (8)], but here it is simplified based on the symbol symmetry shown in Fig. 1, avoiding the need to define five sub-bands as in [10].
@@ V-C Optimum SE Size $N_{\rm e}^{({\rm max\,Capa})}$ for Maximum Rate
V-C1 Existence of an Optimal SE
@@
A close examination of 3GPP BLER results confirms that FDSS-SE performance is non-monotonic with SE size. For example, rate improvements with SE can be observed in [15, Tabs. 2-7], and the achievable-rate formulation in (33) can help to predict such performance behavior in realistic setups. Fig. 12 shows the SNR loss of FDSS-SE relative to plain DFT-s-OFDM versus the SE ratio for $N_{\rm sc}=96$ with 15 kHz subcarrier spacing, and a TDL-C(300 ns) channel. Two FDSS window types are considered: a 3-tap filter equivalent (up to half a subcarrier shift) to a deformed Hann window with $\beta=-11$ dB, and the TRRC window from [10] with parameters $(\rho,\beta)=(0.5,-0.65)$. The theoretical SNR loss is obtained by numerically finding the required SNR – with and without FDSS-SE – in (33) to achieve target spectral efficiencies of $1/3$, $2/3$, and $1$ bpcu. This is compared with 3GPP evaluations of the SNR loss reported in [15, Tabs. 2-7] for reaching $10^{-1}$ BLER with QPSK and code rates $1/6$, $1/3$, and $1/2$. We note that these BLER simulations are not from the present authors, and follows the practical 3GPP assumptions in [15, Tab. 1]. As shown in Fig. 12, the SNR loss also forms a bell-shaped curve as versus the SE ratio, yielding an optimal SE size $N_{\rm e}^{({\rm max\,Capa})}$ from rate-maximization perspective. The theoretical SNR-loss curves exhibit trends consistent with the 3GPP evaluations in [15] – particularly for half-rate coding (1 bpcu) – and can be used to characterize this optimum SE size.
@@
V-C2 Characterizing $N_{\rm e}^{({\rm max\,Capa})}$
Since the combined channel gains $G[k]$ directly depend on the FDSS, the optimal SE size for rate maximization depends on the FDSS shaping. In addition, as $\mathsf{snr}\to\infty$, $g_{0}\approx(1-f(N_{\rm e})/\mathsf{snr})$ for some bounded function $f(N_{\rm e})$ such that the impact of $N_{\rm e}$ on $g_{0}$ diminishes, and $C\approx\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)\log_{2}(\mathsf{snr}/f(N_{\rm e}))$ such that $N_{\rm e}^{({\rm max\,Capa})}\to 0$ as $\mathsf{snr}\to\infty$ in order to maximize the pre-log factor of the capacity. Fig. 13 illustrates $N_{\rm e}^{({\rm max\,Capa})}$ (in % of $N_{\rm sc}$) versus SNR for a Hann FDSS window with $\beta=-11$ dB under both TDL-C (300 ns) delay spread and AWGN channels, with $N_{\rm sc}=96$ and $N_{\rm sc}=600$. The optimum SE size decreases with SNR and tends to zero. Similar to the PAPR case, the optimum ratio $N_{\rm e}^{({\rm max\,Capa})}/N_{\rm sc}$ is nearly bandwidth-invariant. However, under frequency-selective channel, it slightly decreases with larger bandwidth. Overall, $N_{\rm e}^{({\rm max\,Capa})}$ depends primarily on SNR and FDSS shaping.
V-C3 Maximum Rate of FDSS-SE over FDSS
Fig. 14 shows the maximum achievable rate of FDSS-SE when using $N_{\rm e}^{({\rm max\,Capa})}$, compared to FDSS without SE, as a function of FDSS shaping. The channel is TDL-C (300 ns) with $N_{\rm sc}=96$. Both FDSS window types are considered, and plain DFT-s-OFDM corresponds to an FDSS with 0dB attenuation.
V-C4 Rate-PAPR Trade-Off

### v2

@@ IV-C2 Semi-Analytical Characterizations of $N_{\rm e}^{({\rm min\,PAPR})}$ from Bounds
Compared to (15), the bound (13) depends on the complete constellation, as well as $L$. The latter appears thus slightly more precise on Fig. 9 for approximating $N_{\rm e}^{({\rm min\,PAPR})}$. For Fig. 9(e) with 16QAM, we also display the approximate bound (16) which appears to be a slightly better and simpler way to approximate $N_{\rm e}^{({\rm min\,PAPR})}$ than (13) in this case. A defect of these bounds is however that the global optimum SE size obtained from simulations may correspond only to a local minimium in the PAPR bound, as shown in Fig. 9(a) – something we observed for smaller bandwidths and FDSS windows with small attenuation. A simple way to circumvent this issue is to restrict the search range for small FDSS attenuation, or inject a correction factor into the bound. Indeed, we observe that the bounds are further apart for $N_{\rm e}=0$ and become tighter as $N_{\rm e}\to N_{\rm sc}$. A practical method to compensate for this discrepancy is to approximate the PAPR as ${\rm PAPR}\;{\rm[dB]}\approx{\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)\;{\rm[dB]}-K\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)$ where $K$ is the gap between the bound and simulation at $N_{\rm e}=0$, which can empirically be evaluated.
(a) Different QAM orders
@@
Since the search space and summations in  (13) and (15) depend on the bandwidth size, even numerically optimizing these bounds becomes computationally expensive when the bandwidth is large. We therefore consider a further simplification of the bound (15). As illustrated in Fig. 8, once $N_{\rm e}$ increases and the pulses spread apart, the dominant contribution to the peak of the non-coherently combined pulses arises mainly from the overlap of two neighboring pulses, while the sidelobes of the remaining pulses are negligible, especially for larger FDSS attenuation. Accordingly, we reduce the summation in (15) to two pulses as $\max_{n}\left\{\sum_{i,j=0}^{N_{\rm data}-1}|p_{i}[n]||p_{j}[n]|\right\}\approx\max_{n}\left\{\sum_{i,j=0}^{1}|p_{i}[n]||p_{j}[n]|\right\}\!=\!\max_{n}\left\{(|p_{0}[n]|+|p_{1}[n]|)^{2}\right\}$. Since the absolute amplitude level is irrelevant for determining the minimizing location, the square exponent can be omitted. With a small reformulation of the kernel pulse, we thus obtain the approximation
$N_{\rm e}^{({\rm min\,PAPR})}\approx\\ \operatorname*{arg\,min}_{N_{\rm e}}\frac{\max_{n}\left\{w[n]+w\left[n-\frac{1}{N_{\rm sc}-N_{\rm e}}\right]\right\}}{\sqrt{N_{\rm sc}-N_{\rm e}}}$ (20)
where $w[n]=\sqrt{N_{\rm data}}|p_{0}[n]|=\left|\sum_{k=0}^{N_{\rm sc}-1}W[k]e^{{\rm j}\frac{2\pi k}{N_{\rm fft}}n}\right|$ is the magnitude of the $N_{\rm fft}$-IDFT of the zero-padded FDSS window.
Solving (20) analytically appears still intractable, even in the absence of FDSS where the kernel reduces to the closed-form Dirichlet kernel in (8). Nevertheless, (20) provides a simplified connection between the relevant parameters and the PAPR-minimizing SE size, while being computationally very efficient to evaluate.
IV-C3 Comparison with Simulated $N_{\rm e}^{({\rm min\,PAPR})}$
@@
In Fig. 10(b), we show that while the shift value of $L$ can impact the PAPR, the optimal SE size is nearly the same for different $L$ values. For QPSK, $N_{\rm e}^{({\rm min\,PAPR})}$ with $L$ as in (18), (19) and $L=0$ is obtained from simulations, together with the corresponding bound-based estimates. All cases are closely overlapping, and thus the bound (13) does not capture very well the small variation of $N_{\rm e}^{({\rm min\,PAPR})}$ with different $L$ values. In Fig 10(b), $N_{\rm e}^{({\rm min\,PAPR})}$ is also shown for $\pi/2$-BPSK with $L$ as in (18) and $L=0$, from simulation and corresponding bounds. In this case, the range of $N_{\rm e}^{({\rm min\,PAPR})}$ is much smaller and almost constant around $5\%$, indicating that only a small amount of SE may be beneficial. Moreover, contrary to QPSK and other QAMs, as FDSS shaping increases, $N_{\rm e}^{({\rm min\,PAPR})}$ decreases.
In Fig. 11, we compare the simulated $N_{\rm e}^{({\rm min\,PAPR})}$ for different bandwidth sizes $N_{\rm sc}=24,\,96,\,144,\,288,\,480$, and $600$ with the approximation (20). Various configurations of QAM order, frequency shift $L$, and FDSS window are considered. Here, $N_{\rm fft}=8192$, and the simulations target the $10^{-2}$-CCDF PAPR. Both simulations and the approximation in (20) show that, as the bandwidth increases, the value of $N_{\rm e}^{({\rm min\,PAPR})}$ approaches a constant fraction of the bandwidth size. The sawtooth-shaped behavior in (20) stems from the integer constraint on $N_{\rm e}$. As $N_{\rm sc}$ increases, the resolution of $N_{\rm e}/N_{\rm sc}$ improves, and the optimum SE ratio converges to a precise spacing between pulses.
Figure 11: Optimum SE size as function of bandwidth size.
IV-D Simulation Evaluations of Maximum PAPR gain of FDSS-SE over FDSS
Finally, Fig. 12 shows the $10^{-3}$ PAPR CCDF as a function of FDSS shaping, both with and without SE. To reach the lowest PAPR values, the shift parameter $L$ is according to (18) for $\pi/2$-BPSK and (19) for others; and for FDSS-SE, the corresponding optimum SE sizes $N_{\rm e}^{({\rm min\,PAPR})}$ estimated by simulation are used. For QPSK, both filter families – Hann and Kaiser windows – are considered, while for others only results with Hann window are shown for clarity.
For QPSK, SE with Hann window provides an almost constant PAPR gain over FDSS without SE from $1.74$ dB to $1.54$ dB. With Kaiser window, the gain of SE is the same with no FDSS but decreases slightly more to $1.31$ dB. For both window types, the lowest PAPR achieved with FDSS-SE is approximately the same for a given maximum window attenuation ripple. For 16-QAM and 64-QAM, the PAPR gain with Hann window is also constant but smaller, at about $1.3$ and $1.1$ dB, respectively. For $\pi/2$-BSPK, the behavior is different: the gain of SE over FDSS gradually decreases with larger shaping, from 1.51.5dB to 0.30.3dB.
@@ V-A DFT-s-OFDM Reception
where $\mathsf{snr}$ denotes the signal-to-noise ratio (SNR), $\{z[n]\}$ is an additive white Gaussian noise (AWGN) with $\mathbb{E}\left[|z[n]|^{2}\right]=1$, and the channel is normalized so that $\sum_{l=0}^{P-1}\mathbb{E}\left[|h_{p}|^{2}\right]=1$.
@@ V-A1 Equalization
The combining in (25) is equivalent to [31, Eq. (8)], but here it is simplified based on the symbol symmetry shown in Fig. 2, avoiding the need to define five sub-bands as in [31].
@@ V-C Optimum SE Size $N_{\rm e}^{({\rm max\,Capa})}$ for Maximum Rate
V-C1 Observations on the Existence of an Optimal SE
Intuitive Explanation
@@ V-C1 Observations on the Existence of an Optimal SE
BLER Simulations
A close examination of 3GPP BLER results confirms that FDSS-SE performance is non-monotonic with SE size. For example, rate improvements with SE can be observed in [9, Tabs. 2-7], and the achievable-rate formulation in (34) can help to predict such performance behavior in realistic setups. Fig. 14 shows the SNR loss of FDSS-SE relative to plain DFT-s-OFDM versus the SE ratio for $N_{\rm sc}=96$ with 15 kHz subcarrier spacing, and a TDL-C(300 ns) channel. Two FDSS window types are considered: a 3-tap filter equivalent (up to half a subcarrier shift) to a deformed Hann window with $\beta=-11$ dB, and the TRRC window from [31] with parameters $(\rho,\beta)=(0.5,-0.65)$. The theoretical SNR loss is obtained by numerically finding the required SNR – with and without FDSS-SE – in (34) to achieve target spectral efficiencies of $1/3$, $2/3$, and $1$ bpcu. This is compared with 3GPP evaluations of the SNR loss reported in [9, Tabs. 2-7] for reaching $10^{-1}$ BLER with QPSK and code rates $1/6$, $1/3$, and $1/2$. We note that these BLER simulations are not from the present authors, and follows the practical 3GPP assumptions in [9, Tab. 1]. As shown in Fig. 14, the SNR loss evaluated from 3GPP BLER evaluations in [9] and the theoretical SNR loss from (34) form bell-shaped curves versus the SE ratio with consistent trends, particularly for half-rate coding (1 bpcu); such that, the theoretical optimal SE size $N_{\rm e}^{({\rm max\,Capa})}$ matches well the optimal SE size evaluated through BLER simulations.
V-C2 Semi-Analytical Characterization of $N_{\rm e}^{({\rm max\,Capa})}$ from AWGN Capacity
Finding the optimum SE size maximizing the capacity (34) is considerably simpler than optimizing BLER through full link-level simulations. Nevertheless, its evaluation still requires Monte-Carlo averaging to account for the effects of fading channels.
In general, the optimal SE size for rate maximization depends primarily on the FDSS shaping, since the fluctuation in the combined channel gains $G[k]$ mainly depend on the FDSS. To isolate and highlight this dependency, we consider the no-fading AWGN channel case, for which the optimal SE size can be formulated as
$N_{\rm e}^{({\rm max\,AWGN})}=\operatorname*{arg\,max}_{N_{\rm e}}\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)\log_{2}\left(\frac{1}{1-g(N_{\rm e})}\right)$ (35)
where $g(N_{\rm e})=\frac{1}{N_{\rm sc}-N_{\rm e}}\!\!\!\!\!\!\displaystyle\sum_{k=0}^{N_{\rm sc}-N_{\rm e}-1}\!\!\!\!\!\!\frac{1}{1+(\mathsf{snr}W^{c}_{N_{\rm e}}[k])^{-1}}$ with $W^{c}_{N_{\rm e}}[k]=|W[k]|^{2}+|W[k+N_{\rm sc}-N_{\rm e}]|^{2}$ for $0\leq k<N_{\rm e}$ and $W^{c}_{N_{\rm e}}[k]=|W[k]|^{2}$ otherwise.
The expression in (35) can be efficiently evaluated numerically and reveals several expected behaviors of $N_{\rm e}^{({\rm max\,Capa})}$. As $\mathsf{snr}\to\infty$, $g(N_{\rm e})\approx(1-f(N_{\rm e})/\mathsf{snr})$ for some bounded function $f(N_{\rm e})$, such that the variation of $g(N_{\rm e})$ with respect to $N_{\rm e}$ diminishes. Consequently, the AWGN capacity behaves as $C^{\rm AWGN}\approx\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)\log_{2}(\mathsf{snr}/f(N_{\rm e}))$ implying that $N_{\rm e}^{({\rm max\,AWGN})}\to 0$ as $\mathsf{snr}\to\infty$ so as to maximize the pre-log multiplexing factor. In the special case of no FDSS, we obtain $g(N_{\rm e})=\frac{1}{1+\mathsf{snr}^{-1}}\left(N_{\rm sc}-\frac{2N_{\rm e}}{2+\mathsf{snr}^{-1}}\right)$ which is a linearly decreasing function of $N_{\rm e}$, leading again to $N_{\rm e}^{({\rm max\,AWGN})}=0$. Therefore, it can be generally expected that $N_{\rm e}^{({\rm max\,Capa})}$ decreases and tends to zero as SNR increases or as FDSS shaping diminishes. Overall, $N_{\rm e}^{({\rm max\,Capa})}$ depends primarily on SNR and FDSS shaping.
(a) $\beta=11$dB, $\mathsf{snr}=5$dB, 30ns
(b) $\beta=11$dB, $\mathsf{snr}=5$dB, 300ns
(c) $\kappa=2$, $\mathsf{snr}=0$dB, 300ns
(d) $\beta=11$dB, $\mathsf{snr}=0$dB, 300ns
Figure 15: Capacities as a function of SE ratio $N_{\rm e}/N_{\rm sc}$ for different bandwidths $N_{\rm sc}$, Hann ($\beta$) or Kaiser ($\kappa$) FDSS window, $\mathsf{snr}=0$ or 5 dB, and TDL-C channel with 30ns or 300ns delay spread.
@@ V-C2 Semi-Analytical Characterization of $N_{\rm e}^{({\rm max\,Capa})}$ from AWGN Capacity
V-C3 Comparison with Simulated $N_{\rm e}^{({\rm max\,Capa})}$
Fig. 15 shows the achievable rates as a function of SE ratio $N_{\rm e}/N_{\rm sc}$ for different bandwidth sizes $N_{\rm sc}=24,\,144,\,288,\,480$ and $600$. Four configurations are considered. Three cases (Figs. 15(a), 15(b) and 15(d)) use a Hann FDSS window ($\beta=-11$ dB), with $\mathsf{snr}=0$ or $5$ dB, and a TDL-C channel with either 30ns or 300ns delay spread. The remaining case considers a Kaiser window ($\kappa=2$) with $\mathsf{snr}=0$ dB and TDL-C with 300ns delay spread. AWGN capacities for $N_{\rm sc}=24$ and $600$ are shown as references. The AWGN capacity barely changes with different bandwidth size, and only a minor deviation is observed for the Kaiser window. Accordingly, the AWGN rate-optimized SE $N_{\rm e}^{({\rm max\,AWGN})}$ in (35) corresponds to a nearly bandwidth-invariant fraction of the bandwidth. For fading channel with small frequency-selectivity, as in Fig. 15(a), the capacity, and hence the optimal ratio $N_{\rm e}^{({\rm max\,Capa})}/N_{\rm sc}$, remains nearly bandwidth-invariant, similar to the PAPR optimization case. However, under stronger frequency selectivity (Figs. 15(b)-15(d)), the capacity decreases with increasing bandwidth, and the location of the optimal SE exhibits moderate fluctuations. When the bandwidth is small, $N_{\rm e}^{({\rm max\,Capa})}$ is close to $N_{\rm e}^{({\rm max\,AWGN})}$. As the bandwidth increases, the capacity degrades faster without SE than with 50% SE, until a point beyond which the trend reverses. As a result, $N_{\rm e}^{({\rm max\,Capa})}$ temporarily increases, deviates from $N_{\rm e}^{({\rm max\,AWGN})}$, and then converges back toward it. We attribute this behavior to a trade-off between the frequency-diversity gain enabled by SE, which helps mitigate frequency-selective fading, and, the faster convergence of $g_{0}$ toward its mean value as $N_{\rm sc}$ increases, since the summation $g_{0}$ involves a larger number of terms when $N_{\rm e}$ is smaller. Overall, $N_{\rm e}^{({\rm max\,Capa})}$ remains in the vicinity of $N_{\rm e}^{({\rm max\,AWGN})}$, and the capacity curves are relatively flat around their maxima, implying that the capacity obtained with $N_{\rm e}^{({\rm max\,AWGN})}$ and $N_{\rm e}^{({\rm max\,Capa})}$ are very similar.
Fig. 16 shows $N_{\rm e}^{({\rm max\,Capa})}$ (in % of $N_{\rm sc}$) versus SNR for a Hann FDSS window with $\beta=-11$ dB under both AWGN and TDL-C(300ns) channels, for $N_{\rm sc}=96$ and $N_{\rm sc}=600$. In all cases, the optimal SE size decreases with SNR and tends to toward zero, reflecting a fundamental trade-off between frequency-diversity gain and multiplexing efficiency. Edge subcarriers in the excess band experience larger FDSS attenuation and thus operate at lower effective SNR, where frequency diversity from SE is most beneficial. At low SNR, a large number of subcarriers benefit from SE frequency diversity, whereas at high SNR, this diversity gain has less effect and a smaller SE size is preferred to preserve multiplexing gain.
V-D Simulation Evaluations of the Maximum Rate: FDSS-SE versus FDSS
Fig. 17 compares capacities as a function of SNR for an Hann FDSS window with $\beta=-11$ dB. The channel model is TDL-C(300ns) with $N_{\rm sc}=96$. Both FDSS or FDSS-SE incur a capacity loss relative to plain DFT-s-OFDM that increases with SNR. At a spectral efficiency of 1 bpcu, the corresponding SNR losses relative to plain DFT-s-OFDM are 1, 1.4, and 1.8 dB for the rate-optimized FDSS-SE with MRC-MMSE receiver [31], FDSS without SE, and FDSS-SE with basic receiver (see Rem. 1), respectively.
Fig. 18 shows the maximum achievable rate of FDSS-SE when using $N_{\rm e}^{({\rm max\,Capa})}$, compared to FDSS without SE, as a function of FDSS shaping. Both FDSS window types are considered, and plain DFT-s-OFDM corresponds to an FDSS with 0dB attenuation.
Figure 17: Capacity of plain DFT-s-OFDM, rate-optimized FDSS-SE, FDSS-SE with basic receiver, and FDSS only. Hann FDSS window with $\beta=-11$ dB.
@@ V-D Simulation Evaluations of the Maximum Rate: FDSS-SE versus FDSS
V-E Rate-PAPR Trade-Off